OpenStudy (anonymous):

Solve: 6x^2 - 13x = 5 a. {0, 13/6} b. {0, -13/6} c. {-1/3, 5/2} d. {1/3, -5/2}

OpenStudy (anonymous):

\[x = \frac{13 \pm \sqrt{169 - 120}}{12}\]

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

ha I made typo it should be 169+120 =289